OpenStudy (anonymous):

someone help plz ill medal + fan A spinner is divided into ten equal segments. The segments are labeled with the letters A, B, E, I, L, M, O, S, T, U. Match each compound event with its probability. Match Term Definition Probability of A and A A) twelve percent Probability of vowel and consonant B) one percent Probability of M and E and T C) zero point one percent Probability of a letter before J and a letter after R D) twenty-five percent

OpenStudy (anonymous):

do you know how to convert fractions to percents

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok so just do that... take the number of letters the total ofcourse and put that as your denominator

OpenStudy (anonymous):

then your going to find how many a's vowels' and others are in the given numbers then put it over what i said earlier was you denominator then just convert into a percent

OpenStudy (campbell_st):

for compound events, multiply the probabilities for each stage so P(A, A) = 1/10 x 1/10 hope it helps

OpenStudy (anonymous):

you get it

OpenStudy (anonymous):

so what would the number would the numerator would be

OpenStudy (anonymous):

is there choices

OpenStudy (anonymous):

a) 12% b)1% c)0.1% d)25%

OpenStudy (campbell_st):

probability = number of favourable outcomes ---------------------------- total number of outcomes so for P(A) there is only 1 favourable....out of 10 then for the compound event you multiply P(A, A) = 1/10 * 1/10

OpenStudy (anonymous):

either a or d i think

OpenStudy (anonymous):

for every answer can you explain for me

OpenStudy (campbell_st):

so after multiplying its so P(A, A) = 1/100 change it to a percentage

OpenStudy (anonymous):

is it 0.1%

OpenStudy (anonymous):

i think so

OpenStudy (campbell_st):

to change a fraction to a percentages... you multiply the fraction by 100

OpenStudy (campbell_st):

so what is 1/100 * 100 =

OpenStudy (anonymous):

how do you

OpenStudy (anonymous):

do the constant and vowel

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

somebody

OpenStudy (campbell_st):

well that is all the possible outcomes P(vowel and consonant) = 5/10 * 5/10